Show that for any A > 1 and any positive k, "Limit" as n approaches infinity "n^k\/A^n=0"


1
Expert's answer
2022-04-29T13:05:36-0400

Choose an arbitrary number "q" satisfying: "1<q<A". It is enough to prove the following inequality: "\frac{n^k}{A^n}\leq\frac{q^n}{A^n}" for sufficiently large "n" and to show that "\frac{q^n}{A^n}\rightarrow0", "n\rightarrow\infty".

Consider the inequality "n^k\leq q^n". Since the function "ln(x)" is increasing, it is enough to show the inequality: "k\,ln(n)\leq n\, ln(q)". Consider the function: "f(x)=x\,ln(q)-k\,ln(x)". For sufficiently large "x" the following inequality holds: "x^{\frac12}>ln(x)" (it follows from the properties of function "x^{\frac12}-ln(x)" and its derivative). Thus, we obtain: "f(x)=x\,ln(q)-k\,ln(x)>x\,ln(q)-k\,x^{\frac12}=x^{\frac12}(x^{\frac12}ln(q)-k)". We receive that for "x>(\frac{k}{ln(q)})^2" "f(x)>0". Thus, "n\,ln(q)-k\,ln(n)>0" for "n>(\frac{k}{ln(q)})^2". Therefore, "n^k\leq q^n". It remains to prove that "b^n\rightarrow0,n\rightarrow\infty" for "0<b<1".Consider the expression "n\,ln(b)". The latter tends to "-\infty" as "n\rightarrow\infty". "b^n=e^{n\,ln(b)}\rightarrow0,n\rightarrow\infty". It remains to put "b=\frac{q}{A}" to complete the proof. We obtained that "\frac{n^k}{A^n}\leq\frac{q^n}{A^n}\rightarrow0,n\rightarrow\infty."